the variable x and y vary directly. Use the given values to write an equation that relatesx and y.
1) x=3, y=15
2) x=4, y=-16
plz i need help asap

Answer by Alan3354(69443)

the variable x and y vary directly. Use the given values to write an equation that relatesx and y.
1) x=3, y=15
2) x=4, y=-16
------------------
2 points are given, (3,15) and (4,-16)
Find the eqn of the straight line that passes thru them.
Slope m = diffy/diffx = -31/1 = -31
Use the slope and either point to find b, the y-intercept
y = mx + b
15 = 3*(-31) + b
b = 108
--> y = -31x + 108
That's all.